DISPERSED LASER INDUCED FLUORESCENCE OF METALLIC MOLECULAR IONS. IDENTIFICATION OF NEW LOW-LYING ELECTRONIC STATES OF TiCl^{+} AND TiF^{+}

\underline{C. FOCSA}^{1} and B. PINCHEMEL, Laboratoire de Physique des Lasers, Atomes et Molécules, UMR CNRS 8523, Centre d’Etudes et de Recherches Lasers et Applications, Université des Sciences et Technologies de Lille, 59655 Villeneuve d’Ascq, France 1 Present address: Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1.

The dispersed laser-induced fluorescence technique has been applied for the first time to metallic molecular ions. The TiCl^{+} and TiF^{+} ions were produced by a high-voltage discharge in helium with traces of TiCl_{4} or TiF_{4}, respectively. A c.w. dye-laser and a grating spectrometer were used to record low-resolution spectra of these species in the visible region. This led to the the observation of new low-lying electronic states of these ions: the C^{3}Π (∼1535 cm^{−1}) state of TiCl^{+}, the B^{3}Δ (∼2040 cm^{−1}) and C^{3}Π (∼2200 cm^{−1}) states of TiF^{+}. The identification of these new states contributes to a better characterization of the first 3000 cm^{−1} of the energy level diagram of these molecules. The experimental position of the C^{3}Π state of TiCl^{+} is in good agreement with theoretical predictions given by a Ligand Field Theory model.^{a} We have extended these calculations to the TiF^{+} isovalent ion, taking advantage of the new experimental data. Both experimental and theoretical new results presented here are expected to help future high-resolution investigations on these species.

aC. Focsa, M. Bencheikh, and L. G. M. Pettersson, J. Phys. B: At. Mol. Opt. Phys. 31, 2857-2869 (1998).